Find 4 consecutive integers such that twice the sum of the first and third is 11 greater than 3 times the second.
Thank you -

Answer by stanbon(75887)

Find 4 consecutive integers such that twice the sum of the first and third is 11 greater than 3 times the second.
--------------------
1st: x-2
2nd: x-1
3rd: x
4th: x+1
----------------
Equation:
2(x-2 + x) = 3(x-1)+11
------
2(2x-2) = 3x-3+11
------
4x-4 = 3x+8
x = 12
-------
1st: x-2 = 10
2nd: 11
3rd: 12
4th: 13
-----------------
Cheers,
Stan H.